OpenStudy (anonymous):

A boy is 8 years old. His father is 32 years older than he is. What percentage of his father's age is the boy's? A boy is 8 years old. His father is 32 years older than he is. What percentage of his father's age is the boy's? @Mathematics

OpenStudy (anonymous):

8/(32+8) = 8/40. 8/40 x 100 = 20%.